OpenStudy (adi3):

Pls help Will medal how to factor out 13x^2 - 27 = 0

OpenStudy (adi3):

@ganeshie8

OpenStudy (adi3):

@Ahsome

OpenStudy (ahsome):

Is the equation \(13x^2-27=0\)?

OpenStudy (adi3):

yes

OpenStudy (adi3):

sorry the equation is 3x^2 - 27 = 0

OpenStudy (adi3):

not 13, it is 3

OpenStudy (ahsome):

OH. That makes WAY more sense :P

OpenStudy (adi3):

yeah, curently i am angry bc i cant get this equation

OpenStudy (ahsome):

Simple enough :) \(3x^2-27=0\) \(3(x^2-9)=0\) \(3(x^2-3^2)=0\) \(3(x+3)(x-3)=0\)

OpenStudy (adi3):

ok thnxs

OpenStudy (ahsome):

No problem @Adi3 :)